# Optimal preconditioning techniques for finite volume approximation of three-dimensional conservative space-fractional diffusion equations

### v1

@@ 4 Numerical experiments
In the tables, ‘PCG($\tau$)/PGMRES($\tau$)’ means the CG/GMRES method with our proposed preconditioner $\mathbf{P}$, while ‘CG/GMRES’ means those without preconditioner. For further comparison, CG/GMRES with two circulant-type preconditioners, T. Chan’s circulant preconditioner chan2007introduction, fu2019finite and Strang’s circulant preconditioner chan2007introduction, fu2019finite, are also used to solve the corresponding systems. More precisely, all Toeplitz matrices in the coefficient matrix of (2.10) are substituted with the two kinds of circulant preconditioners for preconditioning (2.10) when CG/GMRES method is used, which are denoted as ‘PCG(T)/PGMRES(T)’ and ‘PCG(S)/PGMRES(S)’, respectively. In addition, “CPU” denote the CPU time in seconds for solving the related systems; ‘Iter’ stands for the average iteration numbers of different methods. When the CPU time exceeds 3000 seconds, we stop the iteration and represent the results in the tables as $\dagger$”.
@@ 5 Conclusions
In this paper, we propose sine transform-based preconditioning techniques for the CN-FV scheme of 3D conservative SFDEs. Theoretically, we prove that the proposed preconditioner enables the PCG and PGMRES methods to achieve linear convergence rates for symmetric and non-symmetric cases, respectively. Moreover, our preconditioning techniques not only extend the applicability of the sine transform-based preconditioning techniques from the FD method to the FV method, but also significantly enhance the performance of the circulant preconditioners proposed in fu2019stability, fu2019finite, as demonstrated by numerical results. In future work, we will focus on solving high-dimensional conservative SFDEs with variable coefficients.
@@ Acknowledgments
The work of Wei Qu was supported by the research grant 2024KTSCX069 from the Characteristic Innovation Projects of Ordinary Colleges and Universities in Guangdong Province. The work of Siu-Long Lei was supported by the research grants MYRG-GRG2024-00237-FST-UMDF and MYRG-GRG2023-00181-FST-UMDF from University of Macau. The work of Sean Y. Hon was supported in part by NSFC under grant 12401544 and a start-up grant from the Croucher Foundation.

### v2

@@ 4 Numerical experiments
In the tables, ‘PCG($\tau$)/PGMRES($\tau$)’ means the CG/GMRES method with our proposed preconditioner $\mathbf{P}$, while ‘CG/GMRES’ means those without preconditioner. For further comparison, CG/GMRES with two circulant-type preconditioners, T. Chan’s circulant preconditioner Chan and Jin [2007], Fu et al. [2019a] and Strang’s circulant preconditioner Chan and Jin [2007], Fu et al. [2019a], are also used to solve the corresponding systems. More precisely, all Toeplitz matrices in the coefficient matrix of (2.10) are substituted with the two kinds of circulant preconditioners for preconditioning (2.10) when CG/GMRES method is used, which are denoted as ‘PCG(T)/PGMRES(T)’ and ‘PCG(S)/PGMRES(S)’, respectively. In addition, “CPU” denote the CPU time in seconds for solving the related systems; ‘Iter’ stands for the average iteration numbers of different methods. When the CPU time exceeds 3000 seconds, we stop the iteration and represent the results in the tables as $\dagger$”.
@@ 5 Conclusions
In this paper, we propose sine transform-based preconditioning techniques for the CN-FV scheme of 3D conservative SFDEs. Theoretically, we prove that the proposed preconditioner enables the PCG and PGMRES methods to achieve linear convergence rates for symmetric and non-symmetric cases, respectively. Moreover, our preconditioning techniques not only extend the applicability of the sine transform-based preconditioning techniques from the FD method to the FV method, but also significantly enhance the performance of the circulant preconditioners proposed in Fu et al. [2019b, a], as demonstrated by numerical results. In future work, we will focus on solving high-dimensional conservative SFDEs with variable coefficients.
@@ Acknowledgments
The work of Wei Qu was supported by the research grant 2024KTSCX069 from the Characteristic Innovation Projects of Ordinary Colleges and Universities in Guangdong Province. The work of Siu-Long Lei was supported by the research grants MYRG-GRG2024-00237-FST-UMDF and MYRG-GRG2023-00181-FST-UMDF from University of Macau. The work of Sean Y. Hon was supported in part by NSFC under grant 12401544, the Guangdong and Hong Kong Universities “1+1+1” Joint Research Collaboration Scheme (project no. 2025A0505000014), and a start-up grant from the Croucher Foundation.